QUESTION 2 What is the new mass/volume (m/v)% of a KOH solution that is solution to 330 mL? Ci 2% LLİm 18% prepared by diluting 110 mL of a 6% KOH c) 6% (C) 12% 0 19%
Answer #1

Answer

2%

Explanation

Dilution formula is as follows

C1 × V1 = C2 × V2

C1 = Concentration of initial solution

V1 = Volume of Initial solution

C2 = Concentration of final solution

V2 = Volume of final solution

C2 = C1 × V1/V2

= 6% × 110ml / 330ml

= 2%